OpenStudy (anonymous):

Solve for x 2^(2x-4) x 8^(x-2) = 1/16^(x-2)

OpenStudy (anonymous):

\[(2^{2x-4}) x (8^{x-2}) = 1/16^{x-2}\]

OpenStudy (anonymous):

the x between the brackets is suppose to be a multiply sign

OpenStudy (anonymous):

\[2^{\left( 2x-4 \right)}\times 8^{\left( x-2 \right)}=\frac{ 1 }{ 16^{(x-2)} }=\left(16 \right)^{-\left( x-2 \right)}\]

OpenStudy (anonymous):

write \[8=2^3 and~ 16=2^4\]

OpenStudy (anonymous):

so you need to make all the bases the same number like when adding and subtracting exponents?

OpenStudy (anonymous):

\[8^{x-2}=\left( 2^3 \right)^{(x-2)}=2^{3\left( x-2 \right)}=2^{\left( 3x-6 \right)}\]

OpenStudy (anonymous):

similarly for 16 ,then you can simplify.

OpenStudy (anonymous):

ohhh okay, I think I get it. Thank you!

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

This was my resulting answer 5x-10 log2 2 = -4x + 8 log2 2 5x – 10 = -4x + 8 5x + 4x = 8 +10 9x = 18 x = 18/9 x = 2 I was wondering if it was correct